OpenStudy (djkaras):

how many liters of water must be added to 50 liters of a 30% acid solution in order to produce a 20% acid solution

OpenStudy (yttrium):

Let x be the water added

OpenStudy (yttrium):

.30*50 = .20(50 L +x) .30*50 - .20*50 L = .20x .10*50L = .20x 25L = x Hence, 25 liters of water would need to be added

OpenStudy (djkaras):

Thank you soooooooooooo much helped alot fan me and ill fan you!

OpenStudy (djkaras):



OpenStudy (anonymous):

you're welcome :D

OpenStudy (djkaras):

I was talking to yttrium but thanks any ways!!! :)

OpenStudy (djkaras):

Fan me! and ill meadal You!!!! meadl me to

OpenStudy (djkaras):

Why did you un fan me :C